OpenStudy (anonymous):

At the beginning of an experiment there were 120 grams of a radon, after 42 minutes there were only 96 grams of radon remaining. What is the growth/decay rate of radon?

OpenStudy (anonymous):

\[\frac{96}{120}=e^{-k\times 42}\] solve for \(k\)

OpenStudy (anonymous):

@satellite73 not sure how to

OpenStudy (anonymous):

\[.8=e^{-42k}\] \[\ln(.8)=-42k\] \[k=\frac{\ln(.8)}{-42}\]

OpenStudy (anonymous):

@satellite73 so is -42 the final answer? or ln(.8)/-42

OpenStudy (anonymous):

the second one